# Context-Aware Information Transfer via Digital Semantic Communication in UAV-Based Networks

### v1

@@ I Introduction
To address these limitations, semantic communication has recently gained significant attention. Unlike conventional methods that transmit raw bit-level data, semantic communication focuses on conveying task-relevant or intent-level information [9955312]. In the context of UAV-assisted networks, recent approaches have demonstrated substantial bandwidth savings by transmitting only high-level semantic features rather than full-resolution sensor data. For example, the PE-MMSC framework [guo2025perception] fuses hyperspectral and LiDAR semantics onboard UAVs to reduce transmission volume while maintaining classification accuracy under low-SNR conditions. Similarly, VAE-based encoders have been employed to extract latent semantic representations from UAV imagery, enabling efficient transmission of compressed features that preserve semantic similarity [10437643]. These methods significantly reduce bandwidth consumption while ensuring high reconstruction fidelity and robust task performance for downstream applications such as object detection and scene understanding. While these methods reduce bandwidth consumption and maintain high reconstruction fidelity, they face two key limitations. First, they transmit analog semantic features directly, which makes them highly susceptible to channel noise and difficult to integrate with digital hardware, thereby necessitating digital encoding. Second, they lack context-awareness, which is critical for intelligent surveillance tasks. For instance, if the task is to focus on red cars at high resolution in a congested traffic scenario, context-unaware models may indiscriminately extract all scene elements—such as traffic lights, bicycles, and unrelated vehicles—at lower resolutions, thereby compromising task relevance and overall system efficiency.
@@ II-A Digital Semantic Communication
Quantization plays a pivotal role in digital semantic communication (DSC) by bridging continuous semantic representations and discrete digital signals. Various approaches such as scalar quantization, vector quantization (VQ), and non-linear adaptive schemes have been explored to achieve compact and compatible encodings for digital transmission [zhang2024towards]. Among these, VQ is widely adopted due to its effectiveness in converting high-dimensional semantic embeddings into discrete codewords. This reduces transmission overhead and facilitates integration with digital modulation schemes [zhang2024towards]. In [10065571], VQ-DeepSC incorporates multi-scale embedding with hard quantization via nearest-neighbor search, achieving robust performance under noisy conditions.
However, hard quantization introduces non-differentiability, which affects end-to-end learning, a critical requirement in our system where partial joint optimization of the semantic encoder, UAV network parameters, and decoder is done. To address this, we employ a soft-to-hard quantization strategy, which begins with soft assignments and progressively anneals them into discrete representations. This enables differentiable training while preserving the final discretization necessary for digital transmission. A representative work in this direction is [agustsson2017soft], which proposed soft-to-hard vector quantization for compressible deep representations, demonstrating stable training dynamics and strong compression performance. These quantization approaches, however, have not been fully adapted to UAV communication scenarios where bandwidth and energy limitations amplify the need for quantization-aware design.
@@ III-A UAV State and Energy Consumption
where $\alpha_{\mathrm{move}}$ and $\alpha_{\mathrm{hover}}$ are the average propulsion power coefficients (watts) for movement and hovering. These coefficients are evaluated using the rotary-wing UAV power model [mozaffari2019tutorial]:
@@ IV-B MDP Formulation and Algorithm
Truncated Quantile Critic (TQC) Algorithm: For learning, we employ the TQC algorithm for its robust ability to handle continuous action spaces, mitigate overestimation bias via distributional learning and quantile truncation, and stabilize training through critic ensembling and entropy regularization [kuznetsov2023adapting]. The comprehensive training process for our multi-UAV system using TQC is detailed in Algorithm  1. The core of this learning process involves iteratively updating the actor and critic networks based on experiences sampled from a replay buffer $\mathcal{D}$.
@@ V-A Simulation Setup
The simulation environment was developed using MATLAB R2023b, leveraging the Mapping Toolbox [MATLAB_Mapping_Toolbox_Ref], Deep Learning Toolbox [MATLAB_DL_Toolbox_Ref], and Wireless Communication Toolbox [MATLAB_WC_Toolbox_Ref]. The semantic ViT-CNN encoder-decoder model was implemented in TensorFlow/Keras [TensorFlow_Keras_Ref] and integrated into MATLAB through the ‘pyrun‘ interface [MATLAB_pyrun_Ref], enabling dynamic semantic processing during simulation. For training the semantic encoder-decoder, we used a curated subset of the KITTI raw dataset [Geiger2013IJRR], which consists of $1242\times 375$ RGB images depicting real-world road scenes under diverse lighting and environmental conditions. For testing, a disjoint set of traffic images from an online source [Shutterstock_Germany_Road_Traffic] was used. These were resized and normalized to ensure compatibility with the training data distribution. The mission duration is set to $1000$ seconds to capture long-term system behavior, including UAV scheduling convergence, semantic performance degradation, and GU coverage fairness. Key simulation parameters are summarized in Table I.

### v2

@@ I Introduction
To address these limitations, semantic communication has recently gained significant attention. Unlike conventional methods that transmit raw bit-level data, semantic communication focuses on conveying task-relevant or intent-level information [25]. In the context of UAV-assisted networks, recent approaches have demonstrated substantial bandwidth savings by transmitting only high-level semantic features rather than full-resolution sensor data. For example, the PE-MMSC framework [11] fuses hyperspectral and LiDAR semantics onboard UAVs to reduce transmission volume while maintaining classification accuracy under low-SNR conditions. Similarly, VAE-based encoders have been employed to extract latent semantic representations from UAV imagery, enabling efficient transmission of compressed features that preserve semantic similarity [20]. These methods significantly reduce bandwidth consumption while ensuring high reconstruction fidelity and robust task performance for downstream applications such as object detection and scene understanding. While these methods reduce bandwidth consumption and maintain high reconstruction fidelity, they face two key limitations. First, they transmit analog semantic features directly, which makes them highly susceptible to channel noise and difficult to integrate with digital hardware, thereby necessitating digital encoding. Second, they lack context-awareness, which is critical for intelligent surveillance tasks. For instance, if the task is to focus on red cars at high resolution in a congested traffic scenario, context-unaware models may indiscriminately extract all scene elements—such as traffic lights, bicycles, and unrelated vehicles—at lower resolutions, thereby compromising task relevance and overall system efficiency.
@@ II-A Digital Semantic Communication
Quantization plays a pivotal role in digital semantic communication (DSC) by bridging continuous semantic representations and discrete digital signals. Various approaches such as scalar quantization, vector quantization (VQ), and non-linear adaptive schemes have been explored to achieve compact and compatible encodings for digital transmission [26]. Among these, VQ is widely adopted due to its effectiveness in converting high-dimensional semantic embeddings into discrete codewords. This reduces transmission overhead and facilitates integration with digital modulation schemes [26]. In [7], VQ-DeepSC incorporates multi-scale embedding with hard quantization via nearest-neighbor search, achieving robust performance under noisy conditions.
However, hard quantization introduces non-differentiability, which affects end-to-end learning, a critical requirement in our system where partial joint optimization of the semantic encoder, UAV network parameters, and decoder is done. To address this, we employ a soft-to-hard quantization strategy, which begins with soft assignments and progressively anneals them into discrete representations. This enables differentiable training while preserving the final discretization necessary for digital transmission. A representative work in this direction is [2], which proposed soft-to-hard vector quantization for compressible deep representations, demonstrating stable training dynamics and strong compression performance. These quantization approaches, however, have not been fully adapted to UAV communication scenarios where bandwidth and energy limitations amplify the need for quantization-aware design.
@@ III-A UAV State and Energy Consumption
where $\alpha_{\mathrm{move}}$ and $\alpha_{\mathrm{hover}}$ are the average propulsion power coefficients (watts) for movement and hovering. These coefficients are evaluated using the rotary-wing UAV power model [18]:
@@ IV-B MDP Formulation and Algorithm
Truncated Quantile Critic (TQC) Algorithm: For learning, we employ the TQC algorithm for its robust ability to handle continuous action spaces, mitigate overestimation bias via distributional learning and quantile truncation, and stabilize training through critic ensembling and entropy regularization [14]. The comprehensive training process for our multi-UAV system using TQC is detailed in Algorithm  1. The core of this learning process involves iteratively updating the actor and critic networks based on experiences sampled from a replay buffer $\mathcal{D}$.
@@ V-A Simulation Setup
The simulation environment was developed using MATLAB R2023b, leveraging the Mapping Toolbox [17], Deep Learning Toolbox [5], and Wireless Communication Toolbox [23]. The semantic ViT-CNN encoder-decoder model was implemented in TensorFlow/Keras [1] and integrated into MATLAB through the ‘pyrun‘ interface [4], enabling dynamic semantic processing during simulation. For training the semantic encoder-decoder, we used a curated subset of the KITTI raw dataset [9], which consists of $1242\times 375$ RGB images depicting real-world road scenes under diverse lighting and environmental conditions. For testing, a disjoint set of traffic images from an online source [10] was used. These were resized and normalized to ensure compatibility with the training data distribution. The mission duration is set to $1000$ seconds to capture long-term system behavior, including UAV scheduling convergence, semantic performance degradation, and GU coverage fairness. Key simulation parameters are summarized in Table I.